# Low-Bit Integerization of Vision Transformers using Operand Reodering for Efficient Hardware

### v1

@@ I Introduction
In recent years, Transformers have achieved remarkable success in the natural language processing field, primarily due to their attention mechanism, which enables modeling of dependencies irrespective of their distance in the input or output sequence [1]. Inspired by this success, similar concepts have been applied to various computer vision tasks, including image classification and object detection. While Transformer-based models offer strong performance, their success relies on pretraining using large-scale datasets [2]. One notable limitation is that such pretrained models typically have a fixed architecture—meaning the number of parameters and the computational flow must closely match those of the original model. Consequently, many retraining-based model compression techniques cannot fully leverage the advantages of these pretrained architectures.
To address this challenge, several studies [3, 4] reduce model size by lowering the precision of weights. Although these approaches improve storage efficiency, they often require dequantizing the low-bit parameter back to floating-point to match the level for subsequent modules during inference, resulting in no improvement or extra dequantization cost compared to their floating-point counterparts.
@@ II-B Visual Transformer Hardware Implementation
With the success of ViT in computer vision tasks, both research groups and companies have begun developing hardware implementations to enhance its performance. Although there are several implementation focused on general-purpose processing engine (PE) [8], we focus on customized datapath. A customized datapath enables minimizing memory access and it also successfully reduces power and latency in commercial hardware [9]. For ViTs, Nag [10] built systolic arrays for efficient matrix operations, and Huang [11] further proposed a group-systolic array to balance power and area efficiency compared to the classic systolic array.
@@ III-A Integerization Through Operation Reordering
In contrast, Fig. 1(b) shows the datapath of our integerized ViT. Here, the scaling units are repositioned to follow computational unit, enabling linear layer and matrix multiplication to be executed in low-bit precision. While layer normalization, scaling, and softmax remain in floating point, they are computationally less demanding($O(N^{2})$) compared with the linear layer and matrix multiplication ($O(N^{3})$).
@@ IV Hardware Implementation
To validate our approach and demonstrate its effectiveness, we design hardware that implements the integerized self-attention module. The key components and dataflows are shown in Figure 2. The main modules include linear layer, layer normalization (LayerNorm), matrix multiplication, matrix multiplication with embedded softmax, a reversing module, and multiple quantizers. Dataflow between modules is depicted using triangles (low-bit) or circles (full precision), with channel and time information encoded in color and transparency.
@@ IV-A Low-bit Systolic Array for Linear Layer
The quantized linear layer can be expressed as the product of dequantized input () and weight () with the dequantization achieved by multipliers as shown in Figure 1 (a) with individual step size vector $\Delta_{\mathbf{X}}$ and $\Delta_{\mathbf{W}}$. This linear layer can be written formally as
@@
$\begin{split}\mathbf{Y}&=[\mathbf{X_{q}}(\overline{\Delta_{\mathbf{X}}}\mathbf{I})\mathbf{W_{q}}^{T}]diag(\Delta_{\mathbf{W}})+\mathbf{b}diag(\frac{\Delta_{\mathbf{W}}}{\Delta_{\mathbf{W}}})\\ &=[\mathbf{X_{q}}\mathbf{W_{q}}^{T}+\frac{\mathbf{b}}{\overline{\Delta_{\mathbf{X}}}}diag(\frac{1}{\Delta_{\mathbf{W}}})](\overline{\Delta_{\mathbf{X}}})diag(\Delta_{\mathbf{W}})\end{split}$ (2)
This formulation can be achieved, as shown in Figure 1 (b), with a low-bit linear layer $(\mathbf{X_{q}}\mathbf{W_{q}}^{T})$, an equivalent bias term $(\mathbf{b}(\overline{\Delta_{\mathbf{X}}})^{-1}diag(\frac{1}{\Delta_{\mathbf{W}}}))$ and a post-scaling factor $(diag(\Delta_{\mathbf{W}}))$ with $(\overline{\Delta_{\mathbf{X}}})$ canceled by the subsequent LayerNorm.
@@ IV-B Matrix Multiplication with on-PE Softmax
Matrix multiplication in ViTs is used for both the pre-softmax transformation ($\mathbf{QK^{T}}$) and the weighted sum of value ($\mathbf{W}_{attn}\mathbf{V}$). Although both operations rely on matrix multiplication, they differ in implementation based on subsequent processing.
For $(\mathbf{W}_{attn}\mathbf{V})$, since this matrix multiplication result is passed onto a quantizer, it can be performed at lower bit precision by absorbing the input scales for both operands within the quantizer. We implement this matrix multiplication using a systolic array, as shown in Figure 3: Two input matrices are streamed channel-wise, and each PE contains a local low-bit MAC unit. Final results are latched into a scan chain and pushed out sequentially, while the quantizer is realized with a parallel comparator and an adder.
@@
$\begin{split}attn_{i,j}&=softmax(sQ_{(i,:)}K_{(:,j)})\\ &=exp(sQ_{(i,:)}K_{(:,j)})/\Sigma_{j}exp(sQ_{(i,:)}K_{(:,j)})\\ \end{split}$ (3)
To simplify the exponential operation, we can approximate $e^{x}$ using base-2 exponentiation and integer shift
$\begin{split}&exp(sQ_{(i,:)}K_{(:,j)})\\ &=2\wedge(s\times log2(e)(Q_{(i,:)}K_{(:,j)}))\\ &=2\wedge(\lfloor s\times log2(e)(Q_{(i,:)}K_{(:,j)})\rfloor+r)\\ &=2^{r}<<(\lfloor s\times log2(e)(Q_{(i,:)}K_{(:,j)})\rfloor\\ &\approx(r+1)<<(\lfloor s\times log2(e)(Q_{(i,:)}K_{(:,j)})\rfloor\\ \end{split}$ (4)
@@
Figure 4 demonstrates the hardware to realize this function. The output of matrix multiplication (purple) is passed through the scaled exponential unit. While storing results in the scan chain, the factor $\Sigma_{j}exp(sQ_{(i,:)}K_{(:,j)})$ is calculated and propagated to the end of the row. This sum is fed into the quantizer, built from a series of multiplier with boundary values ($-3.5\Delta_{ATTN}$, …, $1.5\Delta_{ATTN}$, $2.5\Delta_{ATTN}$ in 3-b example) as the comparator reference.
@@ IV-C Systolic-compatible LayerNorm
with initial condition $\mu_{0}=0$ and $\sigma^{2}_{0}=0$. Equation 5 can be realized in a systolic dataflow with a $\mu$ row and a $\sigma^{2}$ row of PE. Its result is broadcasted to the comparator array in Figure 5 to perform the complete the pre-LayerNorm quantization.
@@ V-A Datasets and Training Details
Models. Using a model pretrained from a larger dataset has been shown to achieve higher accuracy compared to training from scratch [2]. Therefore, in this work, we utilize DeiT-S [12] pretrained by Facebook AI. This model is distilled from a larger pretrained model on ImageNet-1k, a dataset comprising 1.28 million training images and 1,000 object classes.
@@ V-B Power Analysis
Q Linear I×$\times$×O 24,576 4.87 10.188 0.414
LayerNorm 2$\times$×O 128 0.03 0.598 4.67
delay $N\times O$ 12,672 0.858
K Linear I×$\times$×O 24,576 4.87 10.188 0.414
LayerNorm 2$\times$×O 128 0.03 0.598 4.67
delay $N\times O$ 12,672 0.858
V Linear I×$\times$×O 24,576 4.87 10.399 0.423
reversing $O\times O$ 1.511
@@ V-C Model Comparison
Int-only Parameter OPs (G) Multiplier CIFAR-10 Accuracy

### v2

# Low-Bit Integerization of Vision Transformers using Operand Reordering for Efficient Hardware

@@ I Introduction
In recent years, Transformers have achieved remarkable success in the natural language processing field, primarily due to their attention mechanism, which enables modeling of dependencies irrespective of their distance in the input or output sequence [1]. Inspired by this success, similar concepts have been applied to various computer vision tasks, including image classification [2] and object detection[3]. While Transformer-based models offer strong performance, their success relies on pretraining using large-scale datasets [2]. One notable limitation is that such pretrained models typically have a fixed architecture—meaning the number of parameters and the computational flow must closely match those of the original model. Consequently, many retraining-based model compression techniques cannot fully leverage the advantages of these pretrained architectures.
To address this challenge, several studies [4, 5] reduce model size by lowering the precision of weights in the same model. Although these approaches improve storage efficiency, they often require dequantizing the low-bit parameter back to floating-point to match the level for subsequent modules during inference [4], resulting in no improvement or extra dequantization cost compared to their floating-point counterparts.
@@ II-B Visual Transformer Hardware Implementation
With the success of ViT in computer vision tasks, both research groups and companies have begun developing hardware implementations to enhance its performance. Although there are several implementation focused on general-purpose processing engine (PE) [9], we focus on customized datapath. A customized datapath enables minimizing memory access and it also successfully reduces power and latency in commercial hardware [10, 11]. For ViTs, Nag [12] built systolic arrays for efficient matrix operations, and Huang [13] further proposed a group-systolic array to balance power and area efficiency compared to the classic systolic array.
@@ III-A Integerization Through Operation Reordering
In contrast, Fig. 1(b) shows the datapath of our integerized ViT. Here, the scaling units are repositioned to follow computational unit, enabling linear layer and matrix multiplication to be executed in low-bit precision. While layer normalization, scaling, and softmax remain in full precision, they are computationally less demanding ($O(N^{2})$) compared with the linear layer and matrix multiplication ($O(N^{3})$).
@@ IV Hardware Implementation
To validate our approach and demonstrate its effectiveness, we design hardware that implements the integerized self-attention module. The key components and dataflows are shown in Figure 2. The main modules include linear layer, layer normalization (LayerNorm), matrix multiplication, matrix multiplication with embedded softmax, a reversing module, and multiple quantizers (q). Dataflow between modules is depicted using triangles (low-bit) or circles (full precision), with channel and time information encoded in color and transparency.
@@ IV-A Low-bit Systolic Array for Linear Layer
The quantized linear layer can be expressed as the product of dequantized input ($\mathbf{X_{q}}$}) and weight ($\mathbf{W_{q}}$}) with the dequantization achieved by multipliers as shown in Figure 1 (a) with individual step size vector $\Delta_{\mathbf{X}}$ and $\Delta_{\mathbf{W}}$. This linear layer can be written formally as
@@
$\begin{split}\mathbf{Y}&\approx[\mathbf{X_{q}}(\overline{\Delta_{\mathbf{X}}}\mathbf{I})\mathbf{W_{q}}^{T}]diag(\Delta_{\mathbf{W}})+\mathbf{b}diag(\frac{\Delta_{\mathbf{W}}}{\Delta_{\mathbf{W}}})\\ &=[\mathbf{X_{q}}\mathbf{W_{q}}^{T}+\frac{\mathbf{b}}{\overline{\Delta_{\mathbf{X}}}}diag(\frac{1}{\Delta_{\mathbf{W}}})](\overline{\Delta_{\mathbf{X}}})diag(\Delta_{\mathbf{W}})\end{split}$ (2)
Equation 2 can be achieved, as shown in Figure 1 (b), with a low-bit linear layer $(\mathbf{X_{q}}\mathbf{W_{q}}^{T})$, an equivalent bias term $(\mathbf{b}(\overline{\Delta_{\mathbf{X}}})^{-1}diag(\frac{1}{\Delta_{\mathbf{W}}}))$ and a post-scaling factor $(diag(\Delta_{\mathbf{W}}))$ with $(\overline{\Delta_{\mathbf{X}}})$ canceled by the subsequent LayerNorm.
@@ IV-B Matrix Multiplication with on-PE Softmax
Matrix multiplication in ViTs is used for both the pre-softmax transformation ($\mathbf{QK^{T}}$) and the weighted sum of value ($\mathbf{W}_{attn}\mathbf{V}$). Although both operations rely on matrix multiplication and can be realized in a weight stationary manner [14], they differ in implementation based on subsequent processing.
For $\mathbf{W}_{attn}\mathbf{V}$, since this matrix multiplication result is passed onto a quantizer, it can be performed at lower bit precision by absorbing the input scales for both operands within the quantizer, similar to Equation 2. We implement this matrix multiplication using a systolic array, as shown in Figure 3: Two input matrices are streamed channel-wise, and each PE contains a local low-bit MAC unit. Final results are latched into a scan chain and pushed out sequentially, while the quantizer is realized with a parallel comparator and an adder.
@@
[…]
To simplify the exponential operation, we approximate $e^{x}$ using base-2 exponentiation and integer shift [5]
[…]
@@
Figure 4 demonstrates the hardware to realize this function. The output of matrix multiplication (purple) is passed through the scaled exponential unit. While storing results in the scan chain, the factor $\Sigma_{j}exp(sQ_{(i,:)}K_{(:,j)})$ is calculated and propagated to the end of the row. This sum is fed into the quantizer, built from a series of multiplier with boundary values ($3.5\Delta_{ATTN}$, …, $1.5\Delta_{ATTN}$, $2.5\Delta_{ATTN}$ in 3-b example) as the comparator reference.
@@ IV-C Systolic-compatible LayerNorm
with initial condition $\mu_{0}=0$ and $\sigma^{2}_{0}=0$. Equation 5 can be realized in a systolic dataflow with a $\mu$ row and a $\sigma^{2}$ row of PE attached at the array, with results sent to the comparator array in Figure 5 for the complete pre-LayerNorm quantization.
@@ V-A Datasets and Training Details
Models. Using a model pretrained from a larger dataset has been shown to achieve higher accuracy compared to training from scratch [2]. Therefore, in this work, we utilize DeiT-S [15] pretrained by Meta. This model is distilled from a larger pretrained model on ImageNet-1k, a dataset comprising 1.28 million training images and 1,000 object classes.
@@ V-B Power Analysis
[…]
@@ V-C Model Comparison
Int-only Parameter
OPs
Multiplier
CIFAR-10
Accuracy